OpenStudy (anonymous):

6x over 5x^2 plus 4x over 5x^2

OpenStudy (zzr0ck3r):

(6x)/(5x^2) + (4x)/(5x^2) = (6x+4x)/(5x^2) = (10x)/(5x^2) = (2x)/(x^2) = 2(x^(1-2))= 2x^(-1) = 2/x